One number is 8 times larger than another. The sum of their reciprocals is 3/8. Determine the two numbers.

Answer by Paul(988)

First number=x
Second number = 8x
Recipocols = 1/x and 1/8x
1/x+1/8x=0.375
(x)+(8x)=0.375[(x)(8x)]
9x=3x^2
3x^2-9x
By using the quadratic formula
x = 3
8(3)=24

Hence the two numbers are 3 and 24.
P.S. If you having problems with quadrtic in this eqaution e-mail me back.
Paul.